OpenStudy (anonymous):

How do you do this problem? X-7= -5

OpenStudy (aum):

Add 7 to both sides of the equation.

OpenStudy (anonymous):

x-7= -5 +7 +7 x=2

OpenStudy (aum):

Correct. x = 2

TheSmartOne (thesmartone):

And \(\Huge\color{red}{Welcome~to~Open~Study}\)